### v1

@@ 3.2.2 Numerical results for triplet correlations in a 2D S/AF heterostructure
Figs. 2(c) and (d) represent the triplet correlations induced at the S/AF interfaces with compensated layered antiferromagnet having two different orientations with respect to the interface. If the ferromagnetic layers are oriented along the interface [Fig. 2(c)], the proximity induced triplet correlations are conventional triplet sign-preserving correlations. The only difference from the case of S/F interface is that their amplitude is strongly weakened due to the fact that the correlations, which are produced by subsequent layers of the antiferromagnet, have the opposite sign and thus partially cancel each other. The cancellation is not full because the ferromagnetic layers, which are far from the interface produce weaker triplet correlations, as it should be expected. The most unexpected thing happens when we consider S/AF interface with ferromagnetic layers oriented perpendicular to the interface [Fig. 2(d)]. In this case each of the AF layers aligned with the interface produce the Néel-type checkerboard triplet correlations. Analogously to Fig. 2(c) the correlations produced by subsequent layers also partially cancel each other resulting in much weaker amplitude of the Néel triplet correlations.
@@ 3.2.3 Numerical results for triplet correlations in a 3D S/AF heterostructure
Figs. 3(c) and 3(d) demonstrate triplet correlations obtained at 3D S/AF interfaces with layered $A$-type AFs in two different orientations with respect to the S/AF interface. In Fig. 3(c) the normal to the plane of ferromagnetic layers is along the S/AF interface. It is seen that in contrast to the 2D S/AF interface with layers oriented perpendicular to the interface [Fig. 2(d)] the triplet correlations are not visible at all. This is because the ferromagnetic order in plane of the interface prevents generation of the checkerboard Néel triplet correlations, and the layered $A$-type Néel triplet correlations are negligibly small, as it was discussed above. It is interesting to compare this situation with the case of S/AF interface with $C$-type antiferromagnet oriented such that the normal to the antiferromagnetic layers is perpendicular to the S/AF interface (along the $x$-axis), see Fig. 3(e). Both situations, presented in Figs. 3(c) and (e) correspond to the same magnetic ordering in the $(x,z)$-plane, but differ by the magnetic ordering in plane of the S/AF interface, which is ferromagnetic for Fig. 3(c) and checkerboard antiferromagnetic for Fig. 3(e). In the second case we see that the checkerboard Néel triplet correlations are induced because the checkerboard AF ordering of the interface allows for their generation. However, the overall amplitude of these checkerboard Néel triplet correlations is weaker than for the S/AF interface with $G$-type antiferromagnet presented in Fig. 3(b) in full analogy with the 2D case considered before.
In Fig. 3(d) the normal to the plane of ferromagnetic layers of the $A$-type antiferromagnet is perpendicular to the S/AF interface. In full analogy with the 2D case [see Fig. 2(c)] weak conventional sign preserving triplet correlations are produced here. The periodic spatial structure of their amplitude is due to the inevitable Friedel-like oscillations of the anomalous Green’s function. The same Friedel-like oscillations are also seen on top of the checkerboard Néel triplet superconducting correlations in Fig. 2(e). For other orientations of the $C$-type AF with respect to the interface the situation should be similar to the presented in Fig. 3(c), that is noticeable triplet correlations are not produced because the layered (not checkerboard) antiferromagnetic ordering in plane of the interface prevent their generation.
@@
Figure 4: Averaged over all superconducting sites superconducting order parameter in 3D superconductor/magnet heterostructures sketched in Fig. 3 as a function of the exchange field of the magnet. Different curves correspond to different heterostructures presented in Fig. 3. Black - S/F heterostructure [Fig. 3(a)]; blue - S/AF heterostructure with $G$-type AF [Fig. 3(b)]; green - S/AF heterostructure with $A$-type AF with normal to the layers oriented along the $x$-axis [Fig. 3(d)]; red - S/AF heterostructure with $A$-type AF with normal to the layers oriented along the $z$-axis [Fig. 3(c)]. $\Delta_{0}=0.035t$, $\Omega_{D}=0.5t$, $T=0.005t$, $\mu=0$.
@@
Although not presented in the figures, the suppression of the OP in homogeneous antiferromagnetic superconductors manifest the same general trends. The strongest suppression occurs in ferromagnetic superconductors [7] and in antiferromagnetic superconductors with $G$-type magnetic ordering at half-filling [80]. However, in antiferromagnetic superconductors with layered magnetic ordering the suppression is very weak and the superconductivity can be fully destroyed only by extremely high exchange fields of the order of the Fermi energy, as it was already reported in [110].
@@ 5 Conclusions
In this paper we investigated the types and magnitude of the Néel triplet superconducting correlations, which are induced in homogeneous antiferromagnetic superconductors and S/AF heterostructures with an arbitrary two-sublattice antiferromagnet. The superconductivity suppression resulting from the singlet-triplet conversion is also studied. It is shown that the appearance of the Néel triplet correlations is very sensitive to the type of the antiferromagnet order. In antiferromagnetic superconductors and S/AF heterostructures with $G$-type antiferromagnetic order near half-filling the checkerboard Néel triplet correlations with approximately the same amplitude as conventional triplets in S/F heterostructures are induced. At the same time, if the exchange field of the antiferromagnet is small as compared to the conduction band width, $A$-type or $C$-type Néel triplet correlations of noticeable amplitude are never generated in antiferromagnetic superconductors and S/AF heterostructures with layered superconductors of $A$ and $C$-type. The amplitude of such correlations is proportional to the small factor $h/t$. The general reason of such discrepancy is the impossibility to reach a good nesting condition in layered superconductors. The strong weakening of the $A$-type or $C$-type Néel triplet correlations at reasonable values of the exchange field with respect to the checkerboard Néel triplet correlations leads to the fact that the conventional singlet superconductivity is suppressed by the layered antiferromagnetic order to a much lesser extent than by the checkerboard antiferromagnetic order at the same value of the exchange field.
@@
Since the triplet correlations is the cornerstone of the majority of effects of superconducting spintronics, the presented study of the superconducting proximity effect and the Néel triplet correlations in S/AF heterostructures is of fundamental importance for an efficient integration of antiferromagnetic materials into the field of superconducting spintronics.
@@ Acknowledgements
G.A.B. and I.V.B. acknowledge the support from Theoretical Physics and Mathematics Advancement Foundation “BASIS” via the project No. 23-1-1-51-1. The analytical analysis was supported by the Russian Science Foundation via the project No. 24-12-00152.

### v2

@@ 3.2.2 Numerical results for triplet correlations in a 2D S/AF heterostructure
Figs. 2(c) and (d) represent the triplet correlations induced at the S/AF interfaces with compensated layered antiferromagnet having two different orientations with respect to the interface. If the ferromagnetic layers are oriented along the interface [Fig. 2(c)], the proximity induced triplet correlations are conventional triplet sign-preserving correlations. The only difference from the case of S/F interface is that their amplitude is strongly weakened due to the fact that the correlations, which are produced by subsequent layers of the antiferromagnet, have the opposite sign and thus partially cancel each other. The cancellation is not full because the ferromagnetic layers, which are far from the interface, produce weaker triplet correlations, as it should be expected. The most unexpected thing happens when we consider S/AF interface with ferromagnetic layers oriented perpendicular to the interface [Fig. 2(d)]. In this case each of the AF layers aligned with the interface produce the Néel-type checkerboard triplet correlations. Analogously to Fig. 2(c) the correlations produced by subsequent layers also partially cancel each other resulting in much weaker amplitude of the Néel triplet correlations.
@@ 3.2.3 Numerical results for triplet correlations in a 3D S/AF heterostructure
Figs. 3(c) and 3(d) demonstrate triplet correlations obtained at 3D S/AF interfaces with layered $A$-type AFs in two different orientations with respect to the S/AF interface. In Fig. 3(c) the normal to the plane of ferromagnetic layers is along the S/AF interface. It is seen that in contrast to the 2D S/AF interface with layers oriented perpendicular to the interface [Fig. 2(d)] the triplet correlations are not visible at all. This is because the layered AF order in plane of the interface prevents generation of the checkerboard Néel triplet correlations, and the layered $A$-type Néel triplet correlations are negligibly small, as it was discussed above. It is interesting to compare this situation with the case of S/AF interface with $C$-type antiferromagnet oriented such that the normal to the antiferromagnetic layers is perpendicular to the S/AF interface (along the $x$-axis), see Fig. 3(e). Both situations, presented in Figs. 3(c) and (e) correspond to the same magnetic ordering in the $(x,z)$-plane, but differ by the magnetic ordering in plane of the S/AF interface, which is layered for Fig. 3(c) and checkerboard antiferromagnetic for Fig. 3(e). In the second case we see that the checkerboard Néel triplet correlations are induced because the checkerboard AF ordering of the interface allows for their generation. However, the overall amplitude of these checkerboard Néel triplet correlations is weaker than for the S/AF interface with $G$-type antiferromagnet presented in Fig. 3(b) in full analogy with the 2D case considered before.
In Fig. 3(d) the normal to the plane of ferromagnetic layers of the $A$-type antiferromagnet is perpendicular to the S/AF interface. In full analogy with the 2D case [see Fig. 2(c)] weak conventional sign preserving triplet correlations are produced here. The periodic spatial structure of their amplitude is due to the inevitable Friedel-like oscillations of the anomalous Green’s function. The same Friedel-like oscillations are also seen on top of the checkerboard Néel triplet superconducting correlations in Fig. 2(e). For other orientations of the $C$-type AF with respect to the interface the situation should be similar to the presented in Fig. 3(c), that is noticeable triplet correlations are not produced because the layered (not checkerboard) antiferromagnetic ordering in plane of the interface prevents their generation.
@@
Figure 4: Superconducting order parameter as a function of the exchange field of the magnet. The value of the OP is averaged over all superconducting sites. Different curves correspond to different 3D superconductor/magnet heterostructures presented in Fig. 3. Black - S/F heterostructure [Fig. 3(a)]; blue - S/AF heterostructure with $G$-type AF [Fig. 3(b)]; green - S/AF heterostructure with $A$-type AF with normal to the layers oriented along the $x$-axis [Fig. 3(d)]; red - S/AF heterostructure with $A$-type AF with normal to the layers oriented along the $z$-axis [Fig. 3(c)]. $\Delta_{0}=0.035t$, $\Omega_{D}=0.5t$, $T=0.005t$, $\mu=0$.
@@
It is worth mentioning that for all curves presented in Fig. 4 the superconducitng OP at zero exchange field of the AF is smaller than the OP of the isolated superconductor, that is $\Delta(h=0)<\Delta_{0}$. The reason is that at $h=0$ the system is not equivalent to an isolated superconducting film with $\Delta=\Delta_{0}$. It is rather equivalent to the superconductor/normal metal heterostructure. The observed suppression of $\Delta$ with respect to $\Delta_{0}$ at $h=0$ is due to the leakage of the superconducting correlations into the N region.
Although not presented in the figures, the suppression of the OP in homogeneous antiferromagnetic superconductors manifest the same general trends. The strongest suppression occurs in ferromagnetic superconductors [7] and in antiferromagnetic superconductors with $G$-type magnetic ordering at half-filling [83]. However, in antiferromagnetic superconductors with layered magnetic ordering the suppression is very weak and the superconductivity can be fully destroyed only by extremely high exchange fields of the order of the Fermi energy, as it was already reported in [115]. The only difference of the antiferromagnetic superconductors from the thin-film S/AF heterostructures is the absence of the leakage of the superconducting correlations into the nonsuperconducting region.
@@ 5 Conclusions
In this paper we investigated the types and magnitude of the Néel triplet superconducting correlations, which are induced in homogeneous antiferromagnetic superconductors and S/AF heterostructures with an arbitrary two-sublattice antiferromagnet. The superconductivity suppression resulting from the singlet-triplet conversion is also studied. It is shown that the appearance of the Néel triplet correlations is very sensitive to the type of the antiferromagnetic order. In antiferromagnetic superconductors and S/AF heterostructures with $G$-type antiferromagnetic order near half-filling the checkerboard Néel triplet correlations with approximately the same amplitude as conventional triplets in S/F heterostructures are induced. At the same time, if the exchange field of the antiferromagnet is small as compared to the conduction band width, $A$-type or $C$-type Néel triplet correlations of noticeable amplitude are never generated in antiferromagnetic superconductors and S/AF heterostructures with layered antiferromagnets of $A$ and $C$-type. The amplitude of such correlations is proportional to the small factor $h/t$. The general reason of such discrepancy is the impossibility to reach a good nesting condition in S/AF heterostructures with antiferromagnets of such types at any parameters of the materials. The strong weakening of the $A$-type or $C$-type Néel triplet correlations at reasonable values of the exchange field with respect to the checkerboard Néel triplet correlations leads to the fact that the conventional singlet superconductivity is suppressed by the layered antiferromagnetic order to a much lesser extent than by the checkerboard antiferromagnetic order at the same value of the exchange field.
@@
Since the triplet correlations is the cornerstone of the majority of effects of superconducting spintronics, the presented study of the superconducting proximity effect and the Néel triplet correlations in S/AF heterostructures is of fundamental importance for an efficient integration of antiferromagnetic materials into the field of superconducting spintronics. In addition, our results indicate that it is better to choose layered rather than checkerboard antiferromagnets for heterostructures if the appearance of triplet correlations is unimportant for the effects under consideration and suppression of superconductivity is undesirable. This situation may arise if the antiferromagnet is used, for example, only to provide an exchange bias for a ferromagnet in some types of planar heterostructures. Another relevant case is if only the electromagnetic interaction between the superconductor and the antiferromagnet is important. Furthermore, our results suggest that the coexistence of antiferromagnetism and $s$-wave superconductivity should rather be characteristic of layered, and not checkerboard antiferromagnets.
@@ Acknowledgements
The financial support from the Russian Science Foundation via the project No.24-22-00186 is acknowledged.